Institute of Astronomy University of Cambridge

Natural Sciences Tripos - Part III Astrophysics Course Guide

2003 – 2004

NST PART III ASTROPHYSICS

Aims and objectives The Part III course is designed to develop students’ analytic, critical and numerical skills to the point that they are equipped to undertake independent research in astrophysics. These aims are advanced through taught courses, which develop students’ mathematical skills through detailed analysis of topics at the forefront of contemporary research, and through the ‘hands on’ experience of a substantial research project. This latter encourages students to develop a critical attitude and an innovative approach to problem solving. This more independent working style is developed under the guidance of a supervisor from the Institute of Astronomy.

Pre-requisites Entry to the course is restricted to those who have taken Option B in Part II Astrophysics, for which students are selected at the end of their second year.

Lecture courses It is recommended that students take four 24-lecture courses (or equivalent) from those offered in Part III of the Mathematical Tripos. At least three of these courses should be selected from a recommended list of courses, compiled annually by the Institute of Astronomy on the grounds of suitability and relevance. The fourth course may be chosen freely from Part III Maths courses and this need not be relevant to astrophysics. The courses offered in Part III Maths vary from year to year and may be found in the Lecture List in The Reporter. From amongst the courses offered in the academic year 2002-2003, the course recommended by the Institute of Astronomy would be: General Relativity, The Structure and Evolution of Stars, Physical Cosmology, Astrophysical Fluid Dynamics, Galaxies, Astrophysical Discs and Magnetohydrodynamics. Usually the majority of courses on the recommended list are taught by members of the Institute of Astronomy. Students should consult the Part III Course Coordinator for guidance about choice of courses. The lecture courses are supported by Examples Classes given by the lecturer. The course material is examined by examinations taken in June.

Research Projects A compulsory element of the course is a substantial research project, extending over two terms, which is undertaken with the guidance of a supervisor from the Institute of Astronomy. The research project accounts for a third of the total marks available for the course. Each year, the Institute produces a list of approved projects, with supervisors, in June of the preceding academic year. Students are recommended to make contact with prospective supervisors at this stage, thus enabling them to undertake some background work over the summer. In October, students must provide 2 project choices to the Project Coordinator by the 2nd Friday of Michaelmas Full Term. Student choice is approved by the Supervisor and Project Coordinator, with project allocation being confirmed by the 3rd Friday of Michaelmas Full Term. The scope of the project extends far beyond a mere literature review and should involve a substantial component of original research. For example, it may involve the reduction, analysis and interpretation of observational data, or else it may be theoretical in nature, perhaps involving a significant amount of computer programming. (Note that the IoA runs a short “Introduction to Unix” course, early in the Michaelmas Term, which will familiarise students with the operating system). The project supervisor offers eight formal supervisions during the project (3 in Michaelmas, 4 in Lent, 1 in Easter). It is the responsibility of the student to make contact with the supervisor and to arrange contact at mutually convenient times. (Note that although supervisors may be willing to offer help outside University Terms, this cannot be guaranteed.) By the end of the Michaelmas Term, students should have prepared a written progress report (maximum 2 sides) for their supervisor and the Project Coordinator. In the last week of Lent Term, students present a 15-minute oral presentation to supervisors and Part III students, this constituting one of the Lent Term supervisions. The final report must be submitted by the 2nd Friday of the Easter Full Term. The report should contain a declaration of originality and a word count (maximum 7500). At a later date in the Easter Term, students are required to deliver a formal (assessed) presentation of 20 minutes to an audience composed of Part III examiners, Part III students and others from the IoA. Each presentation is to be followed by a question and answer session of 10 minutes (chaired by the Project Coordinator). Following the written examinations in June, the examiners reserve the right to examine any candidate viva voce. At this oral examination, the research project, the content of the lecture courses and general issues in astrophysics may be discussed.

Examinations and criteria for marking The Part III examination papers are marked by Assessors appointed by the Department of Mathematics (who are normally the course lecturers) and these marks are relayed to the Part III Astrophysics Examiners who consider them in conjunction with the marks obtained on the project. The Institute of Astronomy Teaching Committee recommends that the NST Part II Astrophysics examiners mark the written examinations and assess their contribution to the overall degree class according to the following criteria: First class marks A candidate placed in the first class will be able to demonstrate a full command and a secure understanding of the examinable material. Scripts will contain substantially correct solutions to most of the quantitative parts of a question, showing a good grasp of mathematical skills. For the essay and questions of an essay nature, first class marks will be awarded for work which is excellent, both in range and in depth of knowledge and in the argument and analysis that it brings to bear. Upper Second class marks, II.1 A candidate placed in the upper second class will be able to demonstrate a good command and some understanding of the examinable material. Scripts will contain solutions to most of the quantitative parts of a question, thereby demonstrating the basic skills involved. For the essay and questions of an essay nature, II.1 marks will be awarded for work that demonstrates knowledge, but which does not provide as impressive a display of understanding, argument and analysis as those in the first class. Lower Second class marks, II.2 A candidate placed in the lower second class will be able to demonstrate some command of the examinable material but with limited understanding. Candidates should demonstrate the ability to make good attempts at the straightforward parts of questions but limited ability to tackle any of the more challenging topics. Answers to questions of a mathematical nature will show an indication of what is required, but fail to proceed sufficiently far into the later parts to demonstrate the skills involved. Essays in this class may often read like prepared material rote learnt for the occasion and fail to be impressive in the range of relevant knowledge and depth of understanding, being superficial in scope or lacking clarity of structure. Third class marks A candidate placed in the third class will be able to demonstrate some knowledge, but have a poor command of the skills expected and very limited understanding of the examinable material. Essays in this class may be unduly brief, lacking in examples or failing to adhere to the rubric, by, for example, answering intelligently, but on material unrelated to the question, or containing some relevant material presented without clear structure or reasoned explanation. Ordinary/Fail A fail mark will be given when a candidate demonstrates little or no knowledge of the material and little or no ability to begin to tackle questions of a mathematical nature. Essays in this class will demonstrate unsatisfactory command of material through a lack of knowledge and an inability to demonstrate any appreciable understanding. It is likely that such answers will be very brief and incomplete, or rambling and irrelevant.

Feedback and Consultation Mechanisms Students are invited to complete feedback questionnaires for each lecture course, which are relayed to the lecturer. In addition, at the end of the year, there is a general feedback questionnaire on the course as a whole, which is considered by the Astrophysics Teaching Committee. Students are however urged to get in contact with the Part III course coordinator at the earliest opportunity if they encounter problems during the year.

Key Contacts

Course Secretary:  Judith Moss  37521  Hoyle Rm 4
Course Coordinator:  Paul Hewett  37507  H19
Teaching Committee Chair:  Jim Pringle  37513  Hoyle Rm 17
IoA Director:  Douglas Gough  37518  Hoyle Rm 49
Director’s Secretary:  Di Sword  37516  Hoyle Rm 48

Students should immediately notify the Course Coordinator if they encounter problems and for general guidance. Day to day queries may be handled by Mrs Judith Moss (Hoyle Building, Room 4).

General Information

It is hoped as a Part III Astrophysics student that you will feel part of the IoA and will participate in some of the activities that maintain its friendly and interactive atmosphere.

Coffee

The IoA staff have coffee (and tea) in the Hoyle building foyer from about 11:00am. Part III Astrophysics students are invited to enjoy coffee with the staff and postgraduate students.

There is no charge for coffee for Part III Astrophysics students, although this may often not be possible given the location and timing of Part III lectures.

Food There are a number of possibilities for lunchtime food provision. Sandwiches and snacks can be obtained on site from the vending machines located in the Hoyle building, whilst on Wednesday the IoA organises a bread and cheese lunch at 12.30 p.m. which precedes the seminars. Off-site there are several possibilities for canteen style food (Churchill College and the Cavendish Laboratory) or cafeteria style food at ‘Wests’, which is located on the other side of Madingley Road, on the access road to the Cavendish Laboratory.

Seminars There are a number of seminars of astronomical interest within various Cambridge departments. Students are encouraged to attend seminars, although the large number of possibilities implies that students need to be selective in those they attend. The IoA. has two regular series of talks: the Wednesday lunchtime talks at 1.15 p.m. (which are preceded by bread and cheese lunch and which usually constitute 2 half-hour talks on specialised research topics) and the Colloquia (Thursdays at 4.30 p.m. during Full Term) which are preceded by tea at 4.00 p.m. and followed by wine at 5.30 p.m. The Colloquia are hour-long talks that generally contain a larger review element, as well as presenting latest scientific results. In addition, the Cavendish Astrophysics Seminar takes place at 4.30 p.m. on Tuesdays while informal lunchtime talks are held at DAMTP (Monday 1 p.m. for the Cambridge Cosmology and Astrophysics Lunch and Tuesday 1 p.m. for Astrophysical Fluid Dynamics and Nonlinear Patterns). All of the above are listed for the forthcoming week on login to the IoA computer system.

Library

The Institute of Astronomy library holds 8000 books and 11000 volumes of astronomical periodicals. Part III students are encouraged to use the library facilities and may sign out books. If you need any help ask the Librarian, Mark Hurn, who has an office in the library area in the Observatory Building.

Photocopying

There is a photocopier in the reprographic room (opposite the vending machines), another outside Judith Moss’s office in the Hoyle building and one in the main library in the Observatory building. Course-related copying is free of charge. For private copying there is a charge of 3p per A4 sheet.

Computing

Part III students are given full access to the IoA Science Cluster. A Common Terminal Are (CTA) is available for student use in the Observatory Building.

Keys

Part III students will be issued with a key (subject to a £5.00 deposit), enabling them out of hours access to the IoA.

Pigeonholes

Part III students are allocated pigeonholes in the Reception area of the Hoyle building which they are encouraged to check regularly for information about the course and about events in the IoA.